OpenStudy (anonymous):

Complete these ordered pairs for this equation. (-1, ), (0, ), (1, ) 2x+y=3 @heather040200

OpenStudy (anonymous):

ok you have the formula 2x+y=3 just plug in the numbers you are given for x and solve

OpenStudy (anonymous):

so 2(-1)+y=3 2(0)+y=3 2(1)+y=3 now just solve

OpenStudy (anonymous):

k